### v1

@@ D.0.1 Validation
Figure D.1: Validation sample-splitting schema. This figure shows how the data is split into training ($m_{i}$), validation ($hpo$) and testing ($p_{i}$) sets for the U.S. and international samples. All models are trained on the U.S. data and then used to generate out-of-sample predictions for both the U.S. and the international sample. Models are retrained each year using an expanding window of data reaching back to 1973. Predictions are generated monthly, always using the most up-to-date model available at the end of each month, e.g. $p_{1}$ predictions represent all monthly predictions during the period of year $1$ and are generated using the model trained on $m_{1}$.
@@ Appendix E Algorithms for approximating probability distribution function and moments from quantiles
Algorithm 1 contains the algorithm for deriving the probability density function from the quantiles. The algorithm is referred to as quantiles-to-density in the paper. Algorithm 2 contains the algorithm for deriving the moments of the probability distribution function. The algorithm is referred to as density-to-moments in the paper, and the algorithm below does not contain the adjustment of the moments described in Section E.3.
@@ E.1 Fallback to linear B-Splines
For larger cap stocks where the density fallback was applied, the cause was manually checked. The density always exhibited a large mass around zero, indicating that the model was predicting a discrete probability of zero return in the following month. Fallback predictions for large cap stocks were often associated with a corporate event, such as being close to official delisting or a recent trading halt. One such case is detailed in Figure E.1, depicting forecasts for October 2014 for Idenix Pharmaceuticals, which was delisted just five days later. Idenix Pharmaceuticals received a tender takeover offer from Merck on June 9, 2014191919The announcement of the acquisition is available here https://www.merck.com/news/merck-to-acquire-idenix/..
@@
2:$gp=100$triangleright Density of grid points
3:$d_{min}=1e-5$triangleright Minimum value for density function
@@
5:$z=0$triangleright Keep only the highest quantile with -1 return
@@
38:d[$\textbf{d}[\textbf{d}<d_{min}]=d_{min}$triangleright Enforce minimum density
39:$w=min(\textbf{d}[(\tilde{\textbf{x}}<=x_{Q10})])$triangleright Enforce monotonicity in the tails
40:$v=argmin(\textbf{d}[(\tilde{\textbf{x}}<=x_{Q10})])$
@@
42:$w=min(\textbf{d}[(\tilde{\textbf{x}}>=x_{Q90})])$
43:$v=argmin(\textbf{d}[(\tilde{\textbf{x}}>=x_{Q90})])$
@@ E.3 Adjustment of moments
The required moment adjustments are estimated using simulated data with known precise theoretical moments. One million non-central Student’s t distributions are generated with various degrees of freedom and noncentrality parameter.  [^20] Non-central Student’s t distribution is defined as $\frac{Y+nc}{\sqrt{V/df}}$ where $Y$ is a standard normal distribution, $V$ is an independent chi-square random variable, $nc$ is noncentrality parameter, and $df$ are degrees of freedom. See scipy.stats.nct documentation for more details on its implementation. Degrees of freedom are randomly selected from {5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 30, 40, 50, 60, 70, 80, 90, 100, 1000, 10000} with replacement and equal probability of each outcome. The noncentrality parameter is randomly selected using uniform distribution of $[-0.5,5]$ interval focusing mostly on positive skewness, as is observed in the empirical equities data. The scale of the generated distributions is set to 0.1 to limit the effects of the fact that stock returns cannot be smaller than $-1$, which is reflected in the density-to-moments algorithm but not in the theoretical distributions. Precise theoretical moments of the distribution are saved, and quantiles of the distribution are fed into the “naive“ density-to-moments algorithm to compute predicted moments. Generated distributions with theoretical kurtosis larger than $20$ are discarded as they would create strong leverage points and skew the estimated parameters to capture these extremes.
@@
$\tilde{v}$ $=v\times(1.0023-0.0021\times s+0.0022\times k_{e})$
$\tilde{s}$ $=0.9950\times s+0.0261\times s^{2}+0.0107\times k_{e}$
$\tilde{k}$ $=3+1.4185\times k_{e}+0.0466\times k_{e}^{2}-0.7395\times s$
Where $v,s,k$ are variance, skewness, and kurtosis from the “naive“ density-to-moments algorithm, $\tilde{v},\tilde{s},\tilde{k}$ are their adjusted counterparts, and $k_{e}$ is excess kurtosis. All of the estimated parameters are significant with t-distribution over $100$ in absolute terms. The fit is high for all the regressions with $R^{2}$ over $0.9$.

### v2

@@ D.1 Validation
Figure D.1: Validation sample-splitting schema. This figure shows how the data is split into training ($m_{i}$), validation ($hpo$) and testing ($p_{i}$) sets for the U.S. and international samples. All models are trained on the U.S. data and then used to generate out-of-sample predictions for both the U.S. and the international sample. Models are retrained each year using expanding window of data reaching back to 1973. Predictions are generated monthly, always using the most up-to-date model available at the end of each month, e.g. $p_{1}$ predictions represent all monthly predictions during the period of year 111 and are generated using the model trained on $m_{1}$.
@@ Appendix E Algorithms for approximating probability distribution function and moments from quantiles
Algorithm 1 contains the algorithm for deriving the probability density function from the quantiles. The algorithm is referred to as quantiles-to-density in the paper. Algorithm 2 contains the algorithm for deriving the moments of the cumulative probaility distribution function. The algorithm is referred to as density-to-moments in the paper, and the algorithm below does not contain the adjustment of the moments described in Section E.3.
@@ E.1 Fallback to linear B-Splines
For larger cap stocks where the density fallback was applied, the cause was manually checked. The density always exhibited a large mass around zero, indicating that the model was predicting a discrete probability of zero return in the following month. Fallback predictions for large cap stocks were often associated with a corporate event, such as being close to official delisting or a recent trading halt. One such case is detailed in Figure E.1, depicting forecasts for October 2014 for Idenix Pharmaceuticals, which was delisted just five days later. Idenix Pharmaceuticals received a tender takeover offer from Merck on June 9, 2014212121The announcement of the acquisition is available here https://www.merck.com/news/merck-to-acquire-idenix/..
@@
[…]
@@
[…]
@@ E.3 Adjustment of moments
The required moment adjustments are estimated using simulated data with known precise theoretical moments. One million non-central Student’s t distributions are generated with various degrees of freedom and noncentrality parameter.  [^22] Non-central Student’s t distribution is defined as $\frac{Y+nc}{\sqrt{V/df}}$ where $Y$ is a standard normal distribution, $V$ is an independent chi-square random variable, $nc$ is noncentrality parameter, and $df$ are degrees of freedom. See scipy.stats.nct documentation for more details on its implementation. Degrees of freedom are randomly selected from {5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 30, 40, 50, 60, 70, 80, 90, 100, 1000, 10000} with replacement and equal probability of each outcome. The noncentrality parameter is randomly selected using uniform distribution of [0.5,5][-0.5,5][ - 0.5 , 5 ] interval focusing mostly on positive skewness, as is observed in the empirical equities data. The scale of the generated distributions is set to 0.1 to limit the effects of the fact that stock returns cannot be smaller than 1-1- 1, which is reflected in the density-to-moments algorithm but not in the theoretical distributions. Precise theoretical moments of the distribution are saved, and quantiles of the distribution are fed into the “naive“ density-to-moments algorithm to compute predicted moments. Generated distributions with theoretical kurtosis larger than 202020 are discarded as they would create strong leverage points and skew the estimated parameters to capture these extremes.
@@
[…]
Where $v,s,k$ are variance, skewness, and kurtosis from the “naive“ density-to-moments algorithm, $\tilde{v},\tilde{s},\tilde{k}$ are their adjusted counterparts, and $k_{e}$ is excess kurtosis. All of the estimated parameters are significant with t-distribution over 100100100 in absolute terms. The fit is high for all the regressions with $R^{2}$ over 0.90.90.9.